OpenStudy (anonymous):

If f(x) = 2x2/(x - 1) for f(-4)

OpenStudy (anonymous):

is the equation like this? \[f(x)=\frac{ 2x^2 }{ x-1}\]

OpenStudy (anonymous):

no Its written like how it was above ^^^

OpenStudy (anonymous):

if so ? the answer is : \[\frac{ 32 }{ -5 }\]

OpenStudy (anonymous):

what does it mean by 2x2 ? i think it is 2x^2

OpenStudy (anonymous):

Yes Its 2^2 I forgot to put that in

OpenStudy (anonymous):

2x^2

OpenStudy (anonymous):

you there ?

OpenStudy (anonymous):

so my answer is correct . means the final answer is : 32/-5

hartnn (hartnn):

to get f(-4) just plug in x=-4 in f(x)

OpenStudy (anonymous):

I did do that but what do u solve for after

hartnn (hartnn):

so, 2 (-4)^2 / (-4-1) =... ?

OpenStudy (anonymous):

12.8 ..??

OpenStudy (anonymous):

is this correct

hartnn (hartnn):

nopes, how u got that ? 2 (-4)^2 / (-4-1) = 2*16/ (-5) = -32/5

OpenStudy (anonymous):

wait no

OpenStudy (anonymous):

Its -6.4 my bad

hartnn (hartnn):

yup, -6.4 is correct